OpenStudy (anonymous):

The sides of a triangle have measures of x, x, and y, and the perimeter of the triangle is 46. The measure y is 4 centimeters longer than the measure x. Write two equations which represent the two sentences. Solve the system you wrote for this using any technique. The sides of a triangle have measures of x, x, and y, and the perimeter of the triangle is 46. The measure y is 4 centimeters longer than the measure x. Write two equations which represent the two sentences. Solve the system you wrote for this using any technique. @Mathematics

OpenStudy (anonymous):

2x+y=46 x+4=y 2x+y=46 x-y=-4 --------- 3x=42 x=14 2(14)+y=46 28+y=46 y=18